OpenStudy (anonymous):

Let T: R^n -> R^n be a linear transformation, and A be the matrix of the linear transformation. Prove that if det(A) does not equal 0 then T is one-to-one ?

OpenStudy (anonymous):

I know that A must be invertible but not sure how to prove it or tie its not the determinant

OpenStudy (anonymous):

*into

OpenStudy (zarkon):

if \(T(x)=T(y)\) then \(Ax=Ay\) and \(Ax-Ay=\vec{0}\) \[A(x-y)=\vec{0}\] if \(det(A)\neq 0\) then \(A\) is invertable and therefore \[x-y=\vec{0}\Rightarrow x=y\] hence \(T\) is one-to-one